OpenStudy (anonymous):

Solve the following equation. 2(h-8)-h=h-16 A.h=8 B.h=-8 C.infinitely many solutions D.no solution WILL GIVE MEDAL FOR CORRECT ANSWER!!!

OpenStudy (anonymous):

2h-16-h=h-16 h-16=h-16 0h=0 0=0 so C